I have the key for this problem but cannot figure out where he gets his values. Need some help. ThanksUse the following information to answer the following 11 questions. Consider a profit maximizing monopoly market where孫90 ans e . 12.00 aand TR-4502.13 OOOQ Consider a profit maximizing monopoly market where:-93-175A2 12,000Q and TR -85QA2 13,0000 (Relevant range of Q is from 1 to 200; Q is in hundreds, you do not have to multiply your answers by 100, but you do have to round your answers to two decimals) 17 What is the profit maximizing Q the firm will produce? 41. 82 18.) What is the profit maximizing P the firm will charge? 19.) A what level of production is ATC minimized? 20.) What is the ATC at the efficient scale? 3493.06 21.) What is the profitx a loss for this firm if it is maximizing profits? tzo7 640.13 22.) If this firm is currently producing 60 units they should increase production False
Answer #1

Answer:

Given

TC=0.9Q^3-175Q^2+12000Q

TR=-85Q^2+13000Q

So Profit P=TR-TC=-85Q^2+13000Q-0.9Q^3+175Q^2-12000Q

P=-0.9Q^3+90Q^2+1000Q

A)

For Maximum Profit dP/dQ=0

-170Q+13000-2.7Q^2+350Q-12000=0

2.7Q^2-180Q-1000=0

So solving for Q=71.82

B) For P=AR

AR=TR/Q=-85Q+13000

So for profit maximising price=-85*71.82+13000=$6895.30

C)

ATC=TC/Q=0.9Q^2-175Q+12000

for Minimum ATC,

dATC/dQ=0

1.8Q-175=0

Q=97.22

D)

ATC at Q=97.22 is

ATC=0.9*(97.22)^2-175*97.22+12000=$3493.06

E)

Maximum Profit for Q=71.82

P=-0.9Q^3+90Q^2+1000Q

P=-0.9*71.82^3+90*71.82^2+1000Q

P=$202640.04

f) True

Since profit maximisation quantity is 71.82 units and we are currently producing 60 units so we need to increase production for maximum profit.